OpenStudy (anonymous):

A={x∈ℤ/|x^2-1|≤35} Find the number of elements of A

OpenStudy (anonymous):

@ganeshie8 @eliassaab @thomaster @aaronq @Abhisar @AccessDenied @Abmon98 @skullpatrol @SolomonZelman @sweetburger @divu.mkr @hartnn @Jemurray3 @mathmale

ganeshie8 (ganeshie8):

|x^2-1| = x^2-1 for x^2>=1 : x^2 -1 <= 35 solve x

OpenStudy (anonymous):

my answer is 13

OpenStudy (anonymous):

is it good?

ganeshie8 (ganeshie8):

Yep ! solution is : [-6, 6] 13 integers is correct, good job !!